Consider the following information Sx=500,Sy=750,Sxy=100 and n=6 the standard error of estimate is


1
Expert's answer
2021-08-10T09:28:55-0400

Solution:

Given, "S_{xx}=500,S_{yy}=750,S_{xy}=100,n=6"

"r=\dfrac{S_{xy}}{\sqrt{S_{xx}S_{yy}}}=\dfrac{100}{\sqrt{500\times750}}=0.1633"

Then, "SS_{residual}=(1-r^2)S_{yy}=(1-0.1633^2)(750)=730"

Now, standard error of estimate"=\sqrt{\dfrac{SS_{residual}}{n-2}}=\sqrt{\dfrac{730}{4}}=13.51"